# Predictive Planner for Autonomous Driving with Consistency Models

### v2

@@ Abstract
Trajectory prediction and planning are essential for autonomous vehicles to navigate safely and efficiently in dynamic environments. Traditional approaches often treat them separately, limiting the ability for interactive planning. While recent diffusion-based generative models have shown promise in multi-agent trajectory generation, their slow sampling is less suitable for high-frequency planning tasks. In this paper, we leverage the consistency model to build a predictive planner that samples from a joint distribution of ego and surrounding agents, conditioned on the ego vehicle’s navigational goal. Trained on real-world human driving datasets, our consistency model generates higher-quality trajectories with fewer sampling steps than standard diffusion models, making it more suitable for real-time deployment. To enforce multiple planning constraints simultaneously on the ego trajectory, a novel online guided sampling approach inspired by the Alternating Direction Method of Multipliers (ADMM) is introduced. Evaluated on the Waymo Open Motion Dataset (WOMD), our method enables proactive behavior such as nudging and yielding, and also demonstrates smoother, safer, and more efficient trajectories and satisfaction of multiple constraints under a limited computational budget.
@@ I Introduction
Traditionally, these two modules usually operate in a decoupled, alternating manner [27, 9], which introduces fundamental limitations. Some approaches integrate data-driven predictions into an optimization-based planner that is inherently reactive rather than proactive [15]. In highly interactive scenarios, such as lane merging where proactive nudging is essential, reactive planning may cause the ego vehicle to be stuck [4, 5]. Also, existing work attempts to incorporate the ego vehicle’s plan into the prediction module [18], but the initial ego plans may already be suboptimal if ignoring future interaction with others. In addition, the alternation between prediction and planning introduces computational inefficiency, making it challenging to meet the demands of real-time operation.
@@
However, diffusion models typically require many sampling steps to obtain high-quality samples, which becomes challenging to meet the demands of real-time operations on autonomous vehicles. To address this limitation, consistency models [40] are designed to generate data directly from noise in a single step or a small number of steps. This significantly improves sample efficiency and maintains high sample quality under limited computational budgets[12, 29].
@@
Experiments show that our consistency model-based planner generates proactive behaviors, e.g., nudging and yielding, through a single forward pass without alternation between prediction and planning. It outperforms both diffusion models and an accelerated variant DDIM [37] with smoother, safer, and more efficient trajectories under limited computational budgets. Compared to transformer-based models, we enable controllable generation that significantly improves satisfaction of multiple trajectory constraints without affecting the original distribution.
@@ II Related Work
Traditionally, trajectory planning is often formulated as a numerical optimization problem, typically decomposed into path planning and speed planning to facilitate real-time operation in dynamic environments [2]. Beyond standalone planning, preliminary effort have integrated trajectory predictions into planner design for collision avoidance [27] and proactive motion planning [3, 33, 38], but either struggle to scale with the number of agents [20] or compromise optimality due to a restrictive exploration of the solution search space [38]. Game-theoretic approaches have been explored to model agent interactions [31], with efficient computation [14] that considers imperfect traffic agents [43], safe actions [42], and uncertainty reduction in the estimates of agent behaviors [17].
@@ III Problem Formulation
At the current timestep $t_{0}\in\mathbb{R}$, the ego agent’s state history over the past $H_{1}\in\mathbb{N}$ timesteps including the current state is given as $S_{e}^{\text{hist}}=(s_{e}^{t_{0}},s_{e}^{t_{0}-1},...,s_{e}^{t_{0}-H_{1}})\in\mathbb{R}^{(H_{1}+1)\times d_{s}}$, where $d_{s}$ is the dimension of the states, e.g. position, velocity, etc. Let $\mathcal{O}$ denote the set of other agents of all types, such as vehicles, pedestrians, and cyclists. Their history states are given for the current timestep $t_{0}$ and the past $H_{1}$ timesteps, denoted by $S_{\mathcal{O}}^{\text{hist}}=(s_{o}^{t_{0}},s_{o}^{t_{0}-1},...,s_{o}^{t_{0}-H_{1}})_{o\in\mathcal{O}}\in\mathbb{R}^{N_{\mathcal{O}}\times(H_{1}+1)\times d_{s}}$, where $N_{\mathcal{O}}=|\mathcal{O}|$. The map information is denoted by $M\in\mathbb{R}^{N_{m}\times c\times d_{m}}$, which includes $N_{m}\in\mathbb{N}$ polylines in the scenario with $c$ points on each polyline and $d_{m}$ attributes per point. We pad and mask the invalid entries. We assume the ego agent’s goal position is given as $s_{e}^{g}\in\mathbb{R}^{d_{s}}$.
Let $\mathcal{O}_{s}\subset\mathcal{O}$ denote the subset of surrounding agents that may interact with the ego agents. Given the information $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M,s_{e}^{g})$, for the future $H_{2}\in\mathbb{N}$ timesteps, we aim to plan the ego agent’s trajectory $S_{e}^{\text{future}}=(s_{e}^{t_{0}+1},s_{e}^{t_{0}+2},...,s_{e}^{t_{0}+H_{2}})\in\mathbb{R}^{H_{2}\times d_{s}}$, and predict the trajectories of surrounding agents $S_{\mathcal{O}_{s}}^{\text{future}}=(s_{o}^{t_{0}+1},s_{o}^{t_{0}+2},...,s_{o}^{t_{0}+H_{2}})_{o\in\mathcal{O}_{s}}\in\mathbb{R}^{N_{\mathcal{O}_{s}}\times(H_{2})\times d_{s}}$, where $N_{\mathcal{O}_{s}}=|\mathcal{O}_{s}|$. This joint prediction and planning naturally captures interactive behaviors between agents. Additionally, the planned trajectory for the ego vehicle should satisfy $N_{c}\in\mathbb{N}$ constraints $c_{i}(S_{e}^{\text{future}})\leq 0,i=1,2,...,N_{c}$, which may include goal reaching, control limits, or other requirements.
@@ IV-A Motion Transformer Encoder
To encode the trajectory history $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$ and map information $M$ as the conditional input for the consistency model, we adopt the encoder architecture from MTR [34, 35]. This transformer-based encoder effectively models scene context and agents’ interactions, and introduces a dense prediction head with a loss function $\mathcal{L}_{\text{encoder}}$ to train this encoder individually. The MTR++ encoder [35] would be suitable for multi-agent settings, but its implementation is not yet available. More importantly, our method is encoder-agnostic, allowing for flexible choice of any suitable scene encoder to be integrated. We represent this encoded feature of the trajectory history and map as $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})\in\mathbb{R}^{(N_{\mathcal{O}}+1)\times D}$ and $\text{MTR}(M)\in\mathbb{R}^{N_{m}\times D}$, respectively, where $D$ is the embedded feature dimension.
@@ IV-B Consistency Model
where $\mathcal{P}[1,T-1]$ is lognormal distribution over the integer set $\{1,2,...,T-1\}$, and $d$ is a pseudo-huber metric function $d(x,y)=\sqrt{||x-y||_{2}^{2}+x^{2}}-c$ [39].
@@ V-A Data pre-processing
We apply a coordinate transformation $\Gamma$ similar to MTR++ [35] that maps each agent’s trajectory into its own local coordinate frame, centered at its position at the current timestep $t_{0}$. The transformed data $(\Gamma(S_{e}^{\text{future}}),\Gamma(S_{\mathcal{O}_{s}}^{\text{future}}))\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times H_{2}\times d_{s}}$ has substantially reduced variance. We then standardize the transformed trajectories to zero mean and unit variance. To preserve the relative spatial relationships among agents, the reference states are collected as $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times d_{s}}$ consisting of each agent’s position at timestep $t_{0}$.
@@ V-B Consistency Model Training
Let $\mathcal{X}$ denote the space of transformed future trajectories $\Gamma(S_{e}^{\text{future}})$ and $\Gamma(S_{\mathcal{O}s}^{\text{future}})$ from Sec. V-A. Let $\mathcal{Y}$ denote the space of conditional inputs, containing encoded historical trajectories $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$, map features $\text{MTR}(M)$ from Sec. IV-A, ego agent’s planning goal $s_{e}^{g}$, and the reference coordinates $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$ from Sec. V-A. Given a sample $y\in\mathcal{Y}$, we leverage the consistency model [40, 39] to sample future trajectories from the conditional probability distribution $p(\cdot|y)$ over $\mathcal{X}$. Each trajectory sample represents one possible joint future behavior of the ego vehicle and surrounding agents.
For training our predictive planner, we jointly train the MTR encoder and the consistency model together with a hybrid loss function, which is a weighted sum of the consistency training loss $\mathcal{L}_{\text{consistency}}$ in Eq. (IV-B) and the dense prediction loss $\mathcal{L}_{\text{encoder}}$ of the MTR encoder from Sec. IV-A as follows:
@@ V-C Planning Constraints Construction
$c_{\text{acc}}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|a^{i}|-a_{\text{limit}},0)$
$c_{\omega}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|\omega^{i}|-\omega_{\text{limit}},0)$ (5)
$c_{\text{goal}}$ enforces goal-reaching for the ego vehicle at the final timestep $t_{0}+H_{2}$. $c_{\text{acc}}$ and $c_{\omega}$ are control limit constraints over the ego agent’s trajectory for a given $a_{\text{limit}}$ and $\omega_{\text{limit}}$. These constraints are inherently in conflict. For example, only moving the trajectory’s final state closer to the goal can increase acceleration violations of the car
@@ V-D Guided Sampling
We find that this alternating update strategy demonstrates effective convergence and empirically leads to lower violations of multiple conflicting constraints, compared to standard guided sampling with vanilla gradient descent in Eq. (6). It also simplifies step size tuning.
@@ VI-A Experiment setup
Method minADE ($m$) minFDE ($m$) Step
@@ VI-A1 Dataset
We evaluate our method on the Waymo Open Motion Dataset (WOMD) [13], which contains 486k traffic scenarios in the training set and 42k scenarios in the interactive validation set to test on. For each agent, including vehicles, pedestrians, and cyclists, the dataset includes 1 second of history, current state, and 8 seconds of future trajectory, sampled at 0.1-second intervals. This sets our history horizon to be $H_{1}=10$ and future horizons to be $H_{2}=80$ timesteps. We identify the surrounding agents with trajectories $S_{\mathcal{O}_{s}}$ by selecting up to 4 agents closest to the ego vehicle based on their trajectory distances, excluding those beyond a 10-meter threshold. This enables us to focus our predictive planner on the most relevant interactive behaviors between the ego vehicle and its immediate surroundings. The goal $s_{e}^{g}$ of the ego agent is chosen as its state at the end of the 8-second future horizon.
@@ VI-B2 Planning Performance under Limited Computational Budget
In Table I, we quantitatively evaluate the realism of the planned trajectory for the ego vehicle through metrics with human groundtruth: Minimum Average Displacement Error (minADE) and Minimum Final Displacement Error (minFDE) [45] over 6 sampled trajectories. The results show that our consistency model, with only 4 sampling steps, captures realistic human driving patterns comparable to Diffusion-10. Moreover, the addition of guided sampling significantly improves the minFDE for the consistency models since it explicitly incorporates goal-reaching constraints.
In Table II, we evaluate the trajectory quality using four metrics 1) angle change as the smoothness of directional transitions, 2) path length as the efficiency of the chosen route, 3) curvature as the sharpness of turns, and 4) collision rate ($2m$ threshold) as safety. Lower values in these metrics indicate smoother, safer, and more efficient trajectories that align with natural driving patterns. Our consistency model achieves superior trajectory quality compared to all Diffusion-based methods, including Diffusion-10, despite its additional sampling steps. While the transformer method has slightly lower angle changes, this is expected given its substantially larger model size (512 vs. our 128 dimensions).
Furthermore, we demonstrate the flexibility of our approach in choosing goal location in Fig. 3. Our consistency model maintains high-quality trajectory planning when targeting novel goal locations not present in the dataset. The current transformer method (MTR) doesn’t support goal selections.
With the same encoder, the computational time for diffusion-based methods directly depends on the number of sampling steps, shown in Table I. Diffusion-4 and DDIM have the same sampling step as our model, but fail to match our performance in trajectory generation across all metrics, as evidenced in Table I and Table II. The guided sampling, introducing a little computational overhead ($\approx 15\%$), provides higher quality trajectory generation and enables controllable planning. Even Diffusion-10, despite its more sampling steps, cannot achieve a comparable performance of trajectory generation to our approach in Table II.

### v3

@@ Abstract
Trajectory prediction and planning are essential for autonomous vehicles to navigate safely and efficiently in dynamic environments. Traditional approaches often treat them separately, limiting the ability for interactive planning. While recent diffusion-based generative models have shown promise in multi-agent trajectory generation, their slow sampling is less suitable for high-frequency planning tasks. In this paper, we leverage the consistency model to build a predictive planner that samples from a joint distribution of ego and surrounding agents, conditioned on the ego vehicle’s navigational goal. Trained on real-world human driving datasets, our consistency model generates higher-quality trajectories with fewer sampling steps than standard diffusion models, making it more suitable for real-time deployment. To enforce multiple planning constraints simultaneously on the ego trajectory, a novel online guided sampling approach inspired by the Alternating Direction Method of Multipliers (ADMM) is introduced. Evaluated on the Waymo Open Motion Dataset (WOMD), our method enables proactive behavior such as nudging and yielding, and also demonstrates smoother, safer, and more efficient trajectories and satisfaction of multiple constraints under a limited computational budget. The project website is at https://anjianli21.github.io/projects/predictive_planner/.
@@ I Introduction
Traditionally, these two modules usually operate in a decoupled, alternating manner [9, 29, 46, 48], which introduces fundamental limitations. Some approaches integrate data-driven predictions into an optimization-based planner that is inherently reactive rather than proactive [16]. In highly interactive scenarios, such as lane merging where proactive nudging is essential, reactive planning may cause the ego vehicle to be stuck [4, 5]. Also, existing work attempts to incorporate the ego vehicle’s plan into the prediction module [19], but the initial ego plans may already be suboptimal if ignoring future interaction with others. In addition, the alternation between prediction and planning introduces computational inefficiency, making it challenging to meet the demands of real-time operation.
@@
However, diffusion models typically require many sampling steps to obtain high-quality samples, which becomes challenging to meet the demands of real-time operations on autonomous vehicles. To address this limitation, consistency models [44] are designed to generate data directly from noise in a single step or a small number of steps. This significantly improves sample efficiency and maintains high quality under limited computational budgets[12, 32].
@@
Experiments show that our consistency model-based planner generates proactive behaviors, e.g., nudging and yielding, without alternation between prediction and planning. It outperforms both diffusion models and an accelerated variant DDIM [41] with smoother, safer, and more efficient trajectories under limited computational budgets. Compared to transformer-based models, we enable controllable generation that significantly improves satisfaction of multiple trajectory constraints without affecting the original distribution.
@@ II Related Work
Traditionally, trajectory planning is often formulated as a numerical optimization problem, typically decomposed into path planning and speed planning to facilitate real-time operation in dynamic environments [2, 49]. Beyond standalone planning, preliminary effort have integrated trajectory predictions into planner design for collision avoidance [29, 47, 22, 30] and proactive motion planning [3, 15, 34, 37], but either struggle to scale with the number of agents [21] or compromise optimality due to a restrictive exploration of the solution search space [42]. Game-theoretic approaches have been explored to model agent interactions [35], with efficient computation [14] that considers imperfect traffic agents [51], safe actions [50], and uncertainty reduction in the estimates of agent behaviors [18].
@@ III Problem Formulation
At the current timestep $t_{0}\in\mathbb{R}$, the ego agent’s state history over the past $H_{1}\in\mathbb{N}$ timesteps including the current state is given as $S_{e}^{\text{hist}}=(s_{e}^{t_{0}},s_{e}^{t_{0}-1},...,s_{e}^{t_{0}-H_{1}})\in\mathbb{R}^{(H_{1}+1)\times d_{s}}$, where $d_{s}$ is the dimension of the states, e.g. position, velocity, etc. Let $\mathcal{O}$ denote the set of other agents of all types, such as vehicles, pedestrians, and cyclists. Their history states are given for the current timestep $t_{0}$ and the past $H_{1}$ timesteps, denoted by $S_{\mathcal{O}}^{\text{hist}}=(s_{o}^{t_{0}},s_{o}^{t_{0}-1},...,s_{o}^{t_{0}-H_{1}})_{o\in\mathcal{O}}\in\mathbb{R}^{N_{\mathcal{O}}\times(H_{1}+1)\times d_{s}}$, where $N_{\mathcal{O}}=|\mathcal{O}|$. The map information is denoted by $M\in\mathbb{R}^{N_{m}\times c_{p}\times d_{m}}$, which includes $N_{m}\in\mathbb{N}$ polylines in the scenario with $c_{p}$ points on each polyline and $d_{m}$ attributes per point. We pad and mask the invalid entries. We assume the ego agent’s goal state is given as $s_{e}^{g}\in\mathbb{R}^{d_{s}}$.
Let $\mathcal{O}_{s}\subset\mathcal{O}$ denote the subset of surrounding agents that may interact with the ego agent. Given the information $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M,s_{e}^{g})$, for the future $H_{2}\in\mathbb{N}$ timesteps, we aim to plan the ego agent’s trajectory $S_{e}^{\text{future}}=(s_{e}^{t_{0}+1},s_{e}^{t_{0}+2},...,s_{e}^{t_{0}+H_{2}})\in\mathbb{R}^{H_{2}\times d_{s}}$, and predict the trajectories of surrounding agents $S_{\mathcal{O}_{s}}^{\text{future}}=(s_{o}^{t_{0}+1},s_{o}^{t_{0}+2},...,s_{o}^{t_{0}+H_{2}})_{o\in\mathcal{O}_{s}}\in\mathbb{R}^{N_{\mathcal{O}_{s}}\times H_{2}\times d_{s}}$, where $N_{\mathcal{O}_{s}}=|\mathcal{O}_{s}|$. This joint prediction and planning naturally captures interactive behaviors between agents. Additionally, the planned trajectory for the ego vehicle should satisfy $N_{c}\in\mathbb{N}$ constraints $c_{i}(S_{e}^{\text{future}})\leq 0,i=1,2,...,N_{c}$, which may include goal reaching, control limits, or other requirements.
@@ IV-A Motion Transformer Encoder
To encode the trajectory history $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$ and map information $M$ as the conditional input for the consistency model, we adopt the encoder architecture from MTR [38, 39]. This transformer-based encoder effectively models scene context and agents’ interactions, and introduces a dense prediction head with a loss function $\mathcal{L}_{\text{encoder}}$ [38] to train this encoder individually. The MTR++ encoder [39] would be suitable for multi-agent settings, but its implementation is not yet available. More importantly, our method is encoder-agnostic, allowing for flexible choice of any suitable scene encoder to be integrated. We represent this encoded feature of the trajectory history and map as $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})\in\mathbb{R}^{(N_{\mathcal{O}}+1)\times D}$ and $\text{MTR}(M)\in\mathbb{R}^{N_{m}\times D}$, respectively, where $D$ is the embedded feature dimension.
@@ IV-B Consistency Model
where $\mathcal{P}[1,T-1]$ is lognormal distribution over the integer set $\{1,2,...,T-1\}$, and $d$ is a pseudo-huber metric function $d(x,y)=\sqrt{||x-y||_{2}^{2}+\delta^{2}}-\delta$ [43] where $\delta$ is a small constant.
@@ V-A Data pre-processing
Instead of using an ego-centric coordinate frame, we apply a coordinate transformation $\Gamma$ similar to MTR++  [39] that maps each agent’s trajectory into its own local coordinate frame, which is centered at each agent’s position at the current timestep $t_{0}$. The transformed data $(\Gamma(S_{e}^{\text{future}}),\Gamma(S_{\mathcal{O}_{s}}^{\text{future}}))\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times H_{2}\times d_{s}}$ has substantially reduced variance. We then standardize the transformed trajectories to zero mean and unit variance. To preserve the relative spatial relationships among agents, the reference states are collected as $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times d_{s}}$ consisting of each agent’s own position at timestep $t_{0}$.
@@ V-B Consistency Model Training
Let $\mathcal{X}$ denote the space of transformed future trajectories $\Gamma(S_{e}^{\text{future}})$ and $\Gamma(S_{\mathcal{O}_{s}}^{\text{future}})$ from Sec. V-A. Let $\mathcal{Y}$ denote the space of conditional inputs, containing encoded historical trajectories $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$, map features $\text{MTR}(M)$ from Sec. IV-A, ego agent’s planning goal $s_{e}^{g}$, and the reference coordinates $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$ from Sec. V-A. Given a sample $y\in\mathcal{Y}$, we leverage the consistency model [44, 43] to sample future trajectories from the conditional probability distribution $p(\cdot|y)$ over $\mathcal{X}$. Each trajectory sample represents one possible joint future behavior of the ego vehicle and surrounding agents.
For training our predictive planner, we jointly train the MTR encoder and the consistency model together with a hybrid loss function, which is a weighted sum of the consistency training loss $\mathcal{L}_{\text{consistency}}$ in Eq. (IV-B) and the dense prediction loss $\mathcal{L}_{\text{encoder}}$ of the MTR encoder from Sec. IV-A with details in [38] as follows:
@@ V-C Planning Constraints Construction
$c_{\text{acc}}=\frac{1}{H_{2}}\sum_{i=t_{0}+1}^{t_{0}+H_{2}}max(|a^{i}|-a_{\text{limit}},0)$
$c_{\omega}=\frac{1}{H_{2}}\sum_{i=t_{0}+1}^{t_{0}+H_{2}}max(|\omega^{i}|-\omega_{\text{limit}},0)$ (5)
$c_{\text{goal}}$ enforces goal-reaching for the ego vehicle at the final timestep $t_{0}+H_{2}$. $c_{\text{acc}}$ and $c_{\omega}$ are control limit constraints over the ego agent’s trajectory for a given $a_{\text{limit}}$ and $\omega_{\text{limit}}$. These constraints are inherently in conflict. For example, only moving the trajectory’s final state closer to the goal can increase acceleration violations of the car.
@@ V-D Guided Sampling
We find that this alternating update strategy demonstrates effective convergence and empirically leads to lower violations of multiple conflicting constraints, compared to standard guided sampling with vanilla gradient descent in Eq. (6). It also simplifies step size tuning. Note that this differs from ADMM since it doesn’t have a dual update step.
@@ VI-A Experiment setup
Method minADE (m) minFDE (m) Sampling Step
MotionDiffuser* [53] 0.533 0.007 32
@@ VI-A1 Dataset
We evaluate our method on the Waymo Open Motion Dataset (WOMD) [13], which contains 486k traffic scenarios in the training set and 42k scenarios in the interactive validation set to test on. For each agent, including vehicles, pedestrians, and cyclists, the dataset includes 1 second of history, current state, and 8 seconds of future trajectory, sampled at 0.1-second intervals. This sets our history horizon to be $H_{1}=10$ and future horizons to be $H_{2}=80$ timesteps. We identify the surrounding agents with trajectories $S_{\mathcal{O}_{s}}$ by selecting up to 4 agents closest to the ego vehicle based on their trajectory distances, excluding those beyond a 10-meter threshold. This choice focuses the predictive planner on the most relevant and interactive neighboring agents, and our model architecture is fully compatible with more agents without additional modifications. The goal $s_{e}^{g}$ of the ego agent is chosen as its state at the end of the 8-second future horizon.
@@ VI-A3 Methods and Baselines
MotionDiffuser[53]
This is state-of-the-art diffusion-based approach for multi-agent trajectory prediction. As the official code has not been released, we report their results of trajectory realism directly from the original paper on the interactive split of the WOMD.
@@ VI-B2 Planning Performance under Limited Computational Budget
In Table I, we quantitatively evaluate the realism of the planned trajectory for the ego vehicle through metrics with human groundtruth: Minimum Average Displacement Error (minADE) and Minimum Final Displacement Error (minFDE) [53] over 6 sampled trajectories. The results show that our consistency model, with only 4 sampling steps, captures realistic human driving patterns comparable to Diffusion-10. Moreover, the addition of guided sampling significantly improves the minFDE for the consistency models since it explicitly incorporates goal-reaching constraints. We also report MotionDiffuser [38] results on joint minADE and minFDE over the pair of interactive agents, which applies an attractor to the GT final point. Our method achieves comparable performance to MotionDiffuser while requiring significantly fewer sampling steps (4 vs. 32).
In Table II, we evaluate the trajectory quality using four metrics 1) Angle change, measuring the smoothness of directional transitions, 2) Path length, measuring the total distance traveled along the planning trajectory to reflect route efficiency, 3) Curvature, quantifying the sharpness of turns, and 4) Collision rate (with a $2m$ threshold), indicating safety. Lower values in these metrics indicate smoother, safer, and more efficient trajectories that align with natural driving patterns. Our consistency model achieves superior trajectory quality compared to all Diffusion-based methods, including Diffusion-10, despite its additional sampling steps. While the transformer method has slightly lower angle changes, this is expected given its substantially larger model size (512 vs. our 128 dimensions).
Furthermore, we demonstrate the flexibility of our approach in choosing goal location in Fig. 3. Our consistency model maintains high-quality trajectory planning when targeting novel goal locations not present in the dataset. The current transformer method doesn’t support goal selections.